A motor exerts torque in a carousel for it to attain a speed of 2.5 rev/s starting


from rest in 3.5 s. Find the work done by the motor if the carousel has a radius


of 27 m and a mass of 1.75 X 105 kg. Consider the carousel to be a thin-walled


cylinder rotating about its center

1
Expert's answer
2021-11-09T10:59:42-0500

The work done is equal to the change in kinetic energy

"W=\frac{I\omega^2}{2}=\frac{mr^2(2\pi n)^2}{4}"

"W=\frac{1.75*10^5*27^2*(2\pi *2.5)^2}{4}=6.2*10^9\:\rm N\cdot m"